How quickly can you do this? Fill appropriate sign. ( ‘’) 
3 / 5 ☐ 2 / 3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3/ 5 = (3 × 3) / (5 × 3)
= 9 / 15
2 / 3 = (2 × 5) / 3 × 5)
= 10 / 15
Here, between 9 / 15 and 10 / 15 both have same denominators. Hence, the fraction having greaternumerator will be the greater.
∴3 / 5 < 2 / 3

Thanks